Status Report for SLAC/DOE IEPM Project

For September & October, 1999

SLAC Home Page
Work completed or in progress since the last report.

Work related to Analysis and Collaborations

PingER Develpoment

New Projects

Several new research assistants (RA's) joined the group this month. There are now five RA's working 20hrs/week and 1 part time person working 5hrs/week. All are Stanford students who come to SLAC between courses. With all this extra manpower we are in a position to move forward on a number of development and analysis projects.


Large day-of-week and hour-of-day delays from SLAC to certain XIWT/IPERF sites were noted by Jeremy Hylton. One hypothesis is that SLAC network traffic is dominated by user traffic, which dies down when users go home at the end of the day, while most of the other networks are dominated by customer usage which is spread more evenly over the day. However, SLAC's connection to ESnet is far from overloaded, and no similar effects are seen to other sites. Another hypothesis is that it is the exchange point at MAE-WEST that suffers the above effect.

Les talked with Vern Paxson on some pathological ping behavior.

Les worked with Chritina Vistoli of INFN/CNAF to fix problems accessing the code. Warren contacted the webmaster at INFN/Napoli to fix a broken reverse traceroute server.

Warren investigated some anomilies in data from the monitoring site at Daresbury where it seemed more packets were being received than were sent. Robin Tasker, the PingER administrator for Daresbury discovered the cause was a bug in ping on Red Hat 6.0. Warren worked with Robin Tasker to install the NIKHEF ping, which doesn't suffer the problem.

Les worked with our friends in Hungary to understand the pathological behavior of Internet performance in Hungary.

Les corresponded with someone at TIFR Bombay to understand why packet loss between ESnet and TIFR was so much better than other U.S. sites to India when the traceroute from SLAC and from TIFR shows the routes to be the same. It is possible there is some kind of shaping involved.

New Sites and New Groups

Warren worked with Shiqi He at HEPNRC to get the new monitoring sites added to the data archiving and plotting and to link into the participants page.

Les worked with Pat Krietz to provide details of the requirements for a possible new monitoring site at BINP in Novosibirsk. Warren responded to an enquiry from NASA about installing a monitoring site.

Warren worked with Dave Millsom and Teresa to create a new Subnet for the IEPM machines.


Warren and Kishan continued to try to understand the loading of the VoIP line with TTCP/UDP. Warren used SNMP on the router and used MRTG to plot the utilization. Warren worked with Becca Nitzan of ESnet to better understand what should happen. Becca reported our findings to Cisco, and in response Becca upgraded the IOS and the RSVP code. In addition there were further routing problems on the testbed.

Les met with Bjorn from NetPredict to discuss using/loaning a Smartbits traffic generator. Bjorn agreed to lend SLAC the machine and Warren worked with Rajiv Krishnamurthy to install a Smart bit box. Rajiv installed the control software and tests have begun.

Warren configured a new machine to sit on the VoIP network next to Nereus to spy on the traffic to/from Nereus. Warren created accounts for Kishan and Rajiv and worked with systems to allow them access to the new machine in case of problems.

Warren and Les worked with Becca Nitzan to Co-ordinate with other users of the testbed

Les looked at the time correlation for the SLAC - LBNL ping RTT data.

Les investigated the Matlab package and created a Fourier transform


Warren worked with Becca Nitzan at ESnet to trouble shoot some issues with the IPv6 set up. Eventually it was found that the solution was that the IOS needed upgrading. Warren worked with SLAC net-admins Tony and Jai-Young to do this.

Warren worked on porting the PingER tools to PingERv6, and porting IPv6 applications to perl using the XS tools.


The project co-ordinators for the RIPE Test traffic project installed a program to generate alarms if the delays break away from a long term average. SLAC has been involved in several alarms. Warren is investigating if this is due to route changes.


Les and Charley worked on analysing the surveyor data to obtain details on loss patterns and defined metrics such as error-free-seconds.

Warren retrieved the data for September, and now also has an account at Advanced to retrieve the data rather than request it from the project co-ordinators each month.


Warren worked with operations and systems to install a new machine called Evagore to serve PingER data from LDAP. Warren created a schema and made the data available. If you have OpenLDAP, you can see a demo by typing

ldapsearch -L -h -p 3900 -b "o=ngi-mon" "(objectclass=*)"

Missing data

There were relatively few glitches this month. SLAC suffered several power cuts. The Monitoring site Stanford failed and the system had to be replaced.


Warren and Les worked with the new R.A.'s to complete the required paperwork and determine the hours they would be at SLAC.


Warren and Les met with Wade Hong during his visit to the HEPiX conference to discuss Canadian Internet connectivity and monitoring tools in use.

Warren and Bjorn met Prof David Cheriton from Stanford and discussed networks and monitoring tools. Prof Cheriton suggested we also talk to Prof Fouad Tobagi and will put us in contact with developers at Cisco.

Les arranged to meet with people at RAL and Daresbury While he will be in the U.K.

Les attended the PPDG meeting at SLAC. His notes are available at

Warren and Les attended a voice conference with XIWT/IPERF.

Work Related to Publicity and Information


Warren, Les and Bill Lidinsky worked on an abstract for CHEP2000 called How No Collisions Make Better Physics. Warren, Les and Becca Nitzan worked on another abstract for CHEP2000 called "1-800-CALL-H.E.P."


Warren presented a talk to the HEPiX/HEPNT conference at SLAC.
Back to Top

Revised 22 October 1999
Comments to